OpenStudy (anonymous):

solve |x - 3| = 2x + 2

OpenStudy (anonymous):

|x-3|-2x=2

OpenStudy (anonymous):

x-3-2x=2

OpenStudy (anonymous):

-x+3-2x=2

OpenStudy (anonymous):

Solve these two.